# mFabric: An Efficient and Scalable Fabric for Mixture-of-Experts Training

### v2

@@ 7.4. Cost Efficiency: Pareto Front Analysis
To better capture the trade-offs between networking cost and performance, we present the Pareto Front analysis of different interconnects in Figure 13. This approach offers a more balanced view that avoids favoring low-cost yet low-performance designs (e.g., TopoOpt or over-subscribed topologies), which may not be practically useful despite their low cost. We observe that mFabric consistently defines the Pareto Front and significantly outperforms Fat-tree and Rail-optimized across all four evaluated models in terms of cost efficiency, which is quantified as a performance-per-dollar metric (inverse of training iteration time normalized by networking cost). At 100 Gbps link bandwidth, mFabric achieves 1.2$\times$ to 1.5$\times$ higher cost-efficiency compared to Fat-tree, with Mixtral 8$\times$7B showing the highest improvement. Moreover, mFabric outperforms Rail-optimized by 1.4$\times$ to 1.5$\times$. At 200 Gbps link bandwidth, At 200 Gbps, the advantage grows to 1.4$\times$ to 1.8$\times$ over Fat-tree and 1.7$\times$ to 1.9$\times$ over Rail-optimized. For 400 Gbps networks, mFabric demonstrates even higher cost-efficiency gains: 2.3$\times$ for Mixtral 8$\times$7B, 2.2$\times$ for Mixtral 8$\times$22B. For DeepSeek-R1, mFabric improves the training cost efficiency by 2.1$\times$ compared to Fat-tree.
Notably, mFabric demonstrates strong cost-efficiency across varying link bandwidths, maintaining a 2.0$\times$–2.4$\times$ advantage over Fat-tree and 2.2$\times$–2.6$\times$ over Rail-optimized even at forward-looking 800 Gbps networks. These gains stem from two factors: 1) mFabric directly connects high-traffic regional GPU pairs with optical circuits, reducing the need for excessive electrical switches and optical transceivers in Fat-tree; 2) unlike Fat-tree’s underutilized uniform bisection bandwidth, mFabric optimizes resource allocation to match MoE’s sparse, non-uniform communication patterns, cutting hardware costs while sustaining high performance.
@@ 7.5. Failure Resiliency
Figure 14. [Simulation] Failure resiliency of mFabric.
@@
NIC failures. Figure 14(a) shows the training performance of mFabric when encountering NIC failures, where indirect forwarding is employed to bypass the failed NIC. We observe that mFabric maintains acceptable performance, with only a 3.3% increase in total training time for the DeepSeek-R1 model. This minor overhead is attributed to the inherent network-wide backup policy, where EPS and OCS provide mutual fallback paths, and the intra-host scale-up domain offers sufficient forwarding bandwidth.
@@
In summary, mFabric exhibits resilience to both network and GPU failures, consistently delivering acceptable performance across the evaluated scenarios.
Figure 15. mFabric with co-packaged optical ports directly attached to GPUs.
@@
8. Look Ahead: High-Radix Scale-Up Domains
So far, we have discussed mFabric as a production-ready system only using commodity OCS (Table 1) and networking equipment (e.g., NICs, transceivers) that are agnostic of local scale-up high-bandwidth domains. In this section, we present a look-ahead study and extend the concept of mFabric to better support the emerging trend of high-radix scale-up domains (e.g., Nvidia GB200 NVL72 system interconnects 72 GPUs in a single scale-up high-bandwidth domain)111111mFabric works with NVL72 by splitting scale-out NICs between OCS and EPS.. In these systems, EP’s all-to-all communications consume more bandwidth in the scale-up high-bandwidth domain than that of scale-out fabrics. Therefore, different from conventional settings, we consider the port of regional OCS to be directly attached to the GPU chip, which is enabled by co-packaged optical I/O (aya, [n. d.]; lig, [n. d.]). Based on this, we envision mFabric evolving towards a regional OCS architecture capable of directly receiving optical signals from xPUs (e.g., GPU, TPU, NPU, etc.), as illustrated in Figure 15. Such an optical switching architecture would enable a long-reach, high-speed (e.g., 4 Tbps or more) regional fabric that sits at the boundary of scale-up and scale-out domains, supporting larger-scale switching fabric compared to state-of-the-art NVL72 copper-based interconnects.
Using packet-level simulations [^12] Same setup as discussed in Section 7.1, we compare the training iteration time of mFabric (with optical I/O) with Nvidia GB200 NVL72 (nvl, [n. d.]a). We consider a 2048-GPU cluster training the state-of-the-art MoE model, DeepSeek-V3 (dee, [n. d.]) (with an EP degree of 128, a PP degree of 16, and a micro-batch size of 240) [^13] Same parameters as in (dee, [n. d.]), with a larger EP size to explore larger training batch sizes and better expert capacities.. The NVL72 system is modeled with a per-GPU NVLink bandwidth of 7.2 Tbps (nvl, [n. d.]a, [n. d.]c) in the scale-up domain, alongside 800 Gbps Ethernet for scale-out communication. We assign 64 GPUs within each NVL72 domain to align with parallelism allocation constraints [^14] In most production practices, only 64 out of 72 GPUs in NVL72 are used because training parallelisms are on the orders of 2 (e.g., 4, 8, 16, 32, etc.).. To ensure a fair comparison, we match mFabric’s total GPU bandwidth in mFabric to NVL72’s 8 Tbps, allocating 800 Gbps to Ethernet and splitting the remaining bandwidth equally between NVLink and mFabric (with optical I/O)’s regional OCS.
Figure 16 presents the normalized training iteration time. We observe that mFabric (with optical I/O) lowers iteration time by 1.3$\times$ compared to NVL72, as it offloads intensive cross-node communication overhead to regional reconfigurable OCS. Yet NVL72 has to use the scale-out networks for cross-node transfers. Additionally, mFabric continues to deliver performance gains even when GPU total I/O bandwidth scales to 16 Tbps in next-generation systems.
@@ 9. Discussion
Additional training parallelisms. Other potential parallelisms for LLM training, such as context parallelism (nvi, [n. d.]a), optimizes the computation of long sequences in the attention module. Notably, its traffic is static and does not overlap with EP’s all-to-all communications. mFabric is ready to accommodate this traffic with its intra-host scale-up fabric or reconfigurable OCS fabric.
@@
Support for model scaling. Currently, MoE models have two scaling trends. On the one hand, recent models like Mixtral (8$\times$7B and 8$\times$22B) and Grok-2 contain a small number of huge experts. This fits into the design of mFabric that contains multiple scale-up networks within each reconfigurable high-bandwidth domain to support huge expert layers. On the other hand, some models (qwe, [n. d.]; DeepSeek-AI et al., 2024; dee, [n. d.]) choose to involve a large number of small experts. Therefore, the radius of the EP all-to-all communications does not grow linearly with the number of experts, as multiple small experts will be packed into one GPU for training efficiency. For example, the state-of-the-art MoE model DeepSeek-V3 (dee, [n. d.]), which uses an EP degree of 64 and a PP degree of 16, can be accommodated within multiple 64-port OCSes in an mFabric fabric.
@@
NIC’s optical fanout options. So far, we consider mFabric’s NIC to be equipped with a dedicated port for each TX/RX channel connecting to the regional reconfigurable OCS. In practice, the mFabric architecture is also compatible with other optical fanout options. For example, bi-directional transceivers using optical circulators merge TX and RX ports into one fiber port (Poutievski et al., 2022; Liu et al., 2023a), and optical breakout cables separate multiple fiber channels within a single MPO/MTP port into individual LC connections (bre, [n. d.]).
@@
Supporting models beyond MoE. While mFabric is primarily optimized for large-scale MoE training, it is also applicable to other non-MoE LLMs, such as GPT-3 (gpt, [n. d.]) and LLaMA (lla, [n. d.]). The core benefit of mFabric is the ability to reconfigure the network topology for non-uniform traffic patterns at a regional radius between the local scale-up networks and global scale-out networks. By blurring the boundaries of scale-up and scale-out networks using reconfigurable OCS, mFabric supports cases where non-MoE LLMs’ non-uniform traffic patterns require optimized network topologies. For example, a ring-based topology serves DP all-reduce traffic in a more cost-efficient way than other topologies. Meanwhile, emerging LLM architectures in the ML community propose to build unconventional MoE-like models on top of dense LLMs that exhibit non-uniform traffic patterns at regional scales (Lin et al., 2024; Sukhbaatar et al., 2024), highlighting the need of mFabric’s reconfigurable network topologies with for distributed training.
@@
Support for Multi-Tenant Training Different from conventional small-scale training jobs, running multiple concurrent MoE training jobs on the same set of GPUs is not preferred in industry practice. For example, Alibaba HPN (Qian et al., 2024) highlights that a single training job already occupies 3K GPUs. Meta shares the experience of training Llama-3 405B model with 15K GPUs (Dubey et al., 2024). DeepSeek-V3 was trained with 2K GPUs (dee, [n. d.]). Moreover, xAI discloses that they train the Grok model with 100K GPUs (gro, [n. d.]b). Nevertheless, if needed, mFabric supports cluster-wide multi-tenancy because its regional OCS high-bandwidth domains can be reconfigured as isolated sub-networks for each small-scale tenant job.
@@ 10. Related work
Reconfigurable networks for distributed training. SiP-ML (Khani et al., 2021) explores silicon photonics for high-bandwidth optical interconnects for optimizing static communications from traditional DP and model-parallel communications. TopoOpt (Wang et al., 2023) proposed optimizing the network topology for distributed training jobs using one-shot optically reconfigurable networks. To the best of our knowledge, mFabric is the first to propose an optically reconfigurable network using commodity hardware for MoE training.
@@
Reconfigurable data center networks. There has been a decades-long research agenda focused on designing reconfigurable networks for data centers (Farrington et al., 2010; Wang et al., 2010; Chen et al., 2012; Porter et al., 2013; Liu et al., 2014; Ballani et al., 2020; Xia et al., 2015; Liu et al., 2015; Chen et al., 2017; Mellette et al., 2017, 2020; Hamedazimi et al., 2014; Wang et al., 2023; Amir et al., 2024; Mellette et al., 2016; Farrington et al., 2013; Saran et al., 2024; Wilson et al., 2024; Amir et al., 2023; Wilson et al., 2023; Amir et al., 2022; Raja et al., 2021; Gerard et al., 2021; Zhang et al., 2021). These proposals target generic data center networks, which are not optimized to provide cost-efficient solutions for large-scale MoE training. In particular, traffic-oblivious solutions, such as RotorNet (Mellette et al., 2017, 2024) and Opera (Mellette et al., 2020), result in suboptimal performance for MoE training, as they cannot deliver timely transfers for bandwidth-intensive all-to-all traffic. Meanwhile, the hardware innovations like Sirius (Ballani et al., 2020) feature faster optical switching latency, hence allowing mFabric to achieve much faster topology reconfiguration delay. Shoal (Shrivastav et al., 2019) proposes reconfigurable electronic circuit switches in rack-scale networks and is not unsuitable for large-scale MoE training.
@@
OCS deployments in data centers. Google has pioneered the deployments of OCS technology in production data centers. Over the years, they have transitioned from electronic packet switching (EPS) to hybrid optical-electrical solutions (Al-Fares et al., 2008; Jain et al., 2013; Singh et al., 2015; Gibson et al., 2022; Poutievski et al., 2022; Liu et al., 2023a; Zu et al., 2024). Early systems like Jupiter (Singh et al., 2015) used Clos topologies with EPS to achieve scalability and high bandwidth but faced limitations in power efficiency and adaptability to dynamic workloads. Subsequent innovations in Jupiter Evolving (Poutievski et al., 2022) introduced OCS to complement EPS in a hybrid architecture that improved cost and power efficiency. Most recently, Lightwave Fabrics from Google (Liu et al., 2023a) used reconfigurable MEMS-based OCS to establish topologies for TPU supercomputers (Jouppi et al., 2023). In particular, it performs one-shot topology reconfiguration prior to training such that the topology remains fixed throughout the training process. In contrast, mFabric proposes runtime topology reconfiguration during training to accommodate dynamic MoE traffic patterns. Furthermore, TPU’s optical interconnect only reconfigures links between 4$\times$4$\times$4 cubes, while the intra-cube topology remains static (3D Torus) during OCS reconfiguration. This fixed intra-cube structure is ill-suited for dynamic all-to-all communication, which requires multi-hop forwarding in response to changing and sparse traffic demands (Jouppi et al., 2023).
@@
OCS for scale-up interconnects. The regionally reconfigurable OCS design in mFabric is applicable to expand the high-bandwidth circuit-switched connectivity enabled by NVLink and NVSwitch. For example, forward-looking techniques like Lightmatter passage optical interconnect (lig, [n. d.]) benefit from mFabric by reconfiguring all GPUs within the same EP group at the chip level. This would enable high-radix on-chip photonic communication with massive bandwidth to handle the communication-intensive demands of both TP and EP in MoE training.
@@
Emerging OCS hardware devices and systems. There are recent proposals on designing novel OCS hardware at server scale for chip-to-chip interconnects (lig, [n. d.]; Seok et al., 2016; Zhang et al., 2022; Bunandar et al., 2024). We would like to highlight that these emerging devices require a system-level design to be practical. Similar to other system-level work (Kumar et al., 2024), mFabric’s regional OCS and its algorithmic designs are compatible with this vibrant line of exploration on novel OCS hardware. We provide further discussions on how mFabric integrates with emerging physical-layer optical technologies in Appendix E.

### v3

@@ 7.4. Cost Efficiency: Pareto Front Analysis
To better capture the trade-offs between networking cost and performance, we present the Pareto Front analysis of different interconnects in Figure 13. This approach offers a more balanced view that avoids favoring low-cost yet low-performance designs (e.g., TopoOpt or over-subscribed topologies), which may not be practically useful despite their low cost. We observe that MixNet consistently defines the Pareto Front and significantly outperforms Fat-tree and Rail-optimized across all four evaluated models in terms of cost efficiency, which is quantified as a performance-per-dollar metric (inverse of training iteration time normalized by networking cost). At 100 Gbps link bandwidth, MixNet achieves 1.2$\times$ to 1.5$\times$ higher cost-efficiency compared to Fat-tree, with Mixtral 8$\times$7B showing the highest improvement. Moreover, MixNet outperforms Rail-optimized by 1.4$\times$ to 1.5$\times$. At 200 Gbps link bandwidth, At 200 Gbps, the advantage grows to 1.4$\times$ to 1.8$\times$ over Fat-tree and 1.7$\times$ to 1.9$\times$ over Rail-optimized. For 400 Gbps networks, MixNet demonstrates even higher cost-efficiency gains: 2.3$\times$ for Mixtral 8$\times$7B, 2.2$\times$ for Mixtral 8$\times$22B. For DeepSeek-R1, MixNet improves the training cost efficiency by 2.1$\times$ compared to Fat-tree.
Notably, MixNet demonstrates strong cost-efficiency across varying link bandwidths, maintaining a 2.0$\times$–2.4$\times$ advantage over Fat-tree and 2.2$\times$–2.6$\times$ over Rail-optimized even at forward-looking 800 Gbps networks. These gains stem from two factors: 1) MixNet directly connects high-traffic regional GPU pairs with optical circuits, reducing the need for excessive electrical switches and optical transceivers in Fat-tree; 2) unlike Fat-tree’s underutilized uniform bisection bandwidth, MixNet optimizes resource allocation to match MoE’s sparse, non-uniform communication patterns, cutting hardware costs while sustaining high performance.
@@ 7.5. Failure Resiliency
Figure 14. [Simulation] Failure resiliency of MixNet.
@@
NIC failures. Figure 14(a) shows the training performance of MixNet when encountering NIC failures, where indirect forwarding is employed to bypass the failed NIC. We observe that MixNet maintains acceptable performance, with only a 3.3% increase in total training time for the DeepSeek-R1 model. This minor overhead is attributed to the inherent network-wide backup policy, where EPS and OCS provide mutual fallback paths, and the intra-host scale-up domain offers sufficient forwarding bandwidth.
@@
In summary, MixNet exhibits resilience to both network and GPU failures, consistently delivering acceptable performance across the evaluated scenarios.
8. Look Ahead: High-Radix Scale-Up Domains
So far, we have discussed MixNet as a production-ready system only using commodity OCS (Table 2) and networking equipment (e.g., NICs, transceivers) that are agnostic of local scale-up high-bandwidth domains. In this section, we present a look-ahead study and extend the concept of MixNet to better support the emerging trend of high-radix scale-up domains (e.g., Nvidia GB200 NVL72 system interconnects 72 GPUs in a single scale-up high-bandwidth domain) [^9] MixNet works with NVL72 by splitting scale-out NICs between OCS and EPS.. In these systems, EP’s all-to-all communications consume more bandwidth in the scale-up high-bandwidth domain than that of scale-out fabrics. Therefore, different from conventional settings, we consider the port of regional OCS to be directly attached to the GPU chip, which is enabled by co-packaged optical I/O (aya, [n. d.]; lig, [n. d.]). Based on this, we envision MixNet evolving towards a regional OCS architecture capable of directly receiving optical signals from xPUs (e.g., GPU, TPU, NPU, etc.), as illustrated in Figure 15. Such an optical switching architecture would enable a long-reach, high-speed (e.g., 4 Tbps or more) regional fabric that sits at the boundary of scale-up and scale-out domains, supporting larger-scale switching fabric compared to state-of-the-art NVL72 copper-based interconnects.
Figure 15. MixNet with co-packaged optical ports directly attached to xPUs.
@@ 8. Look Ahead: High-Radix Scale-Up Domains
Using packet-level simulations [^10] Same setup as discussed in Section 7.1, we compare the training iteration time of MixNet (with optical I/O) with Nvidia GB200 NVL72 (nvl, [n. d.]a). We consider a 2048-GPU cluster training the state-of-the-art MoE model, DeepSeek-V3 (dee, [n. d.]) (with an EP degree of 128, a PP degree of 16, and a micro-batch size of 240) [^11] Same parameters as in (dee, [n. d.]), with a larger EP size to explore larger training batch sizes and better expert capacities.. The NVL72 system is modeled with a per-GPU NVLink bandwidth of 7.2 Tbps (nvl, [n. d.]a, [n. d.]c) in the scale-up domain, alongside 800 Gbps Ethernet for scale-out communication. We assign 64 GPUs within each NVL72 domain to align with parallelism allocation constraints [^12] In most production practices, only 64 out of 72 GPUs in NVL72 are used because training parallelisms are on the orders of 2 (e.g., 4, 8, 16, 32, etc.).. To ensure a fair comparison, we match MixNet’s total GPU bandwidth in MixNet to NVL72’s 8 Tbps, allocating 800 Gbps to Ethernet and splitting the remaining bandwidth equally between NVLink and MixNet (with optical I/O)’s regional OCS.
Figure 16 presents the normalized training iteration time. We observe that MixNet (with optical I/O) lowers iteration time by 1.3$\times$ compared to NVL72, as it offloads intensive cross-node communication overhead to regional reconfigurable OCS. Yet NVL72 has to use the scale-out networks for cross-node transfers. Additionally, MixNet continues to deliver performance gains even when GPU total I/O bandwidth scales to 16 Tbps in next-generation systems.
@@ 9. Discussion
Additional training parallelisms. Other potential parallelisms for LLM training, such as context parallelism (nvi, [n. d.]a), optimizes the computation of long sequences in the attention module. Notably, its traffic is static and does not overlap with EP’s all-to-all communications. MixNet is ready to accommodate this traffic with its intra-host scale-up fabric or reconfigurable OCS fabric.
@@
Support for model scaling. Currently, MoE models have two scaling trends. On the one hand, recent models like Mixtral (8$\times$7B and 8$\times$22B) and Grok-2 contain a small number of huge experts. This fits into the design of MixNet that contains multiple scale-up networks within each reconfigurable high-bandwidth domain to support huge expert layers. On the other hand, some models (qwe, [n. d.]; DeepSeek-AI et al., 2024; dee, [n. d.]) choose to involve a large number of small experts. Therefore, the radius of the EP all-to-all communications does not grow linearly with the number of experts, as multiple small experts will be packed into one GPU for training efficiency. For example, the state-of-the-art MoE model DeepSeek-V3 (dee, [n. d.]), which uses an EP degree of 64 and a PP degree of 16, can be accommodated within multiple 64-port OCSes in an MixNet fabric.
@@
NIC’s optical fanout options. So far, we consider MixNet’s NIC to be equipped with a dedicated port for each TX/RX channel connecting to the regional reconfigurable OCS. In practice, the MixNet architecture is also compatible with other optical fanout options. For example, bi-directional transceivers using optical circulators merge TX and RX ports into one fiber port (Poutievski et al., 2022; Liu et al., 2023a), and optical breakout cables separate multiple fiber channels within a single MPO/MTP port into individual LC connections (bre, [n. d.]).
@@
Supporting models beyond MoE. While MixNet is primarily optimized for large-scale MoE training, it is also applicable to other non-MoE LLMs, such as GPT-3 (gpt, [n. d.]) and LLaMA (lla, [n. d.]). The core benefit of MixNet is the ability to reconfigure the network topology for non-uniform traffic patterns at a regional radius between the local scale-up networks and global scale-out networks. By blurring the boundaries of scale-up and scale-out networks using reconfigurable OCS, MixNet supports cases where non-MoE LLMs’ non-uniform traffic patterns require optimized network topologies. For example, a ring-based topology serves DP all-reduce traffic in a more cost-efficient way than other topologies. Meanwhile, emerging LLM architectures in the ML community propose to build unconventional MoE-like models on top of dense LLMs that exhibit non-uniform traffic patterns at regional scales (Lin et al., 2024; Sukhbaatar et al., 2024), highlighting the need of MixNet’s reconfigurable topologies with for distributed training.
@@
Support for Multi-Tenant Training Different from conventional small-scale training jobs, running multiple concurrent MoE training jobs on the same set of GPUs is not preferred in industry practice. For example, Alibaba HPN (Qian et al., 2024) highlights that a single training job already occupies 3K GPUs. Meta shares the experience of training Llama-3 405B model with 15K GPUs (Dubey et al., 2024). DeepSeek-V3 was trained with 2K GPUs (dee, [n. d.]). Moreover, xAI discloses that they train the Grok model with 100K GPUs (gro, [n. d.]b). Nevertheless, if needed, MixNet supports cluster-wide multi-tenancy because its regional OCS high-bandwidth domains can be reconfigured as isolated sub-networks for each small-scale tenant job.
Comparability with other OCS technologies. MixNet’s regional reconfiguration design is orthogonal to the selection of OCS technologies. We note the emergence of new OCS technologies, such as iPronics (ipr, [n. d.]), which offer microsecond-scale reconfiguration latencies and low per-port costs comparable to MEMS-based OCS. These advancements present a promising opportunity for MixNet to reconfigure the topology for every all-to-all communication phase in MoE training, further improving training performance while maintaining cost efficiency. Furthermore, in the main text, we mainly discuss OCS technologies that can actively switch optical signals at the network core. Other emerging OCS techniques, such as tunable laser at endpoints (e.g., transceivers) with arrayed waveguide grating router (AWGR)-based optical switch that pushes reconfigurability into the network endpoints while leaving the network core to be passive (Ballani et al., 2020; Zhang et al., 2024; Ye et al., 2012), are also compatible with MixNet. For example, assigning dedicated wavelengths to critical expert pairs eliminates bandwidth contention for all-to-all EP communication while organizing ring-based topologies for all-reduced communication of TP. However, these AWGR and endpoint-based OCS are primarily in the research stage and not commercially ready for mass production yet, leaving a longer way to be integrated into today’s computer systems.
@@ 10. Related work
Reconfigurable networks for distributed training. SiP-ML (Khani et al., 2021) explores silicon photonics for high-bandwidth optical interconnects for optimizing static communications from traditional DP and model-parallel communications. TopoOpt (Wang et al., 2023) proposed optimizing the network topology for distributed training jobs using one-shot optically reconfigurable networks. To the best of our knowledge, MixNet is the first to propose an optically reconfigurable network using commodity hardware for MoE training.
@@
Reconfigurable data center networks. There has been a decades-long research agenda focused on designing reconfigurable networks for data centers (Farrington et al., 2010; Wang et al., 2010; Chen et al., 2012; Porter et al., 2013; Liu et al., 2014; Ballani et al., 2020; Xia et al., 2015; Liu et al., 2015; Chen et al., 2017; Mellette et al., 2017, 2020; Hamedazimi et al., 2014; Wang et al., 2023; Amir et al., 2024; Mellette et al., 2016; Farrington et al., 2013; Saran et al., 2024; Wilson et al., 2024; Amir et al., 2023; Wilson et al., 2023; Amir et al., 2022; Raja et al., 2021; Gerard et al., 2021; Zhang et al., 2021; Yu et al., 2017). These proposals target generic data center networks, which are not optimized to provide cost-efficient solutions for large-scale MoE training. In particular, traffic-oblivious solutions, such as RotorNet (Mellette et al., 2017, 2024) and Opera (Mellette et al., 2020), result in suboptimal performance for MoE training, as they cannot deliver timely transfers for bandwidth-intensive all-to-all traffic. Meanwhile, the hardware innovations like Sirius (Ballani et al., 2020) feature faster optical switching latency, hence allowing MixNet to achieve much faster topology reconfiguration delay. Shoal (Shrivastav et al., 2019) proposes reconfigurable electronic circuit switches in rack-scale networks and is not unsuitable for large-scale MoE training.
@@
OCS deployments in data centers. Google has pioneered the deployments of OCS technology in production data centers. Over the years, they have transitioned from electronic packet switching (EPS) to hybrid optical-electrical solutions (Al-Fares et al., 2008; Jain et al., 2013; Singh et al., 2015; Gibson et al., 2022; Poutievski et al., 2022; Liu et al., 2023a; Zu et al., 2024). Early systems like Jupiter (Singh et al., 2015) used Clos topologies with EPS to achieve scalability and high bandwidth but faced limitations in power efficiency and adaptability to dynamic workloads. Subsequent innovations in Jupiter Evolving (Poutievski et al., 2022) introduced OCS to complement EPS in a hybrid architecture that improved cost and power efficiency. Most recently, Lightwave Fabrics from Google (Liu et al., 2023a) used reconfigurable MEMS-based OCS to establish topologies for TPU supercomputers (Jouppi et al., 2023). In particular, it performs one-shot topology reconfiguration prior to training such that the topology remains fixed throughout the training process. In contrast, MixNet proposes runtime topology reconfiguration during training to accommodate dynamic MoE traffic patterns. Furthermore, TPU’s optical interconnect only reconfigures links between 4$\times$4$\times$4 cubes, while the intra-cube topology remains static (3D Torus) during OCS reconfiguration. This fixed intra-cube structure is ill-suited for dynamic all-to-all communication, which requires multi-hop forwarding in response to changing and sparse traffic demands (Jouppi et al., 2023).
@@
OCS for scale-up interconnects. The regionally reconfigurable OCS design in MixNet is applicable to expand the high-bandwidth circuit-switched connectivity enabled by NVLink and NVSwitch. For example, forward-looking techniques like Lightmatter passage optical interconnect (lig, [n. d.]) benefit from MixNet by reconfiguring all GPUs within the same EP group at the chip level. This would enable high-radix on-chip photonic communication with massive bandwidth to handle the communication-intensive demands of both TP and EP in MoE training.
@@
Emerging OCS hardware devices and systems. There are recent proposals on designing novel OCS hardware at server scale for chip-to-chip interconnects (lig, [n. d.]; Seok et al., 2016; Zhang et al., 2022; Bunandar et al., 2024). We would like to highlight that these emerging devices require a system-level design to be practical. Similar to other system-level work (Kumar et al., 2024), MixNet’s regional OCS and its algorithmic designs are compatible with this vibrant line of exploration on novel OCS hardware.